OpenStudy (anonymous):

Use integration by parts to solve: integral ln(x^2) dx

OpenStudy (anonymous):

\[\int\limits \ln(x^2)dx\]

OpenStudy (kainui):

Any ideas?

OpenStudy (anonymous):

I am just not sure which one to use as the derivative and the antiderivative.

OpenStudy (kainui):

It looks like you don't have any choice here, you can only possibly choose: \[\LARGE u=\ln(x^2) \ \ \ \ \ \ dv=dx\]

OpenStudy (anonymous):

then use chain rule on \[\ln (x^2)\]